# Monotonic Paraphrasing Improves Generalization of Language Model Prompting

### v1

@@ Abstract
Performance of large language models (LLMs) may vary with different prompts or instructions of even the same task. One commonly recognized factor for this phenomenon is the model’s familiarity with the given prompt or instruction, which is typically estimated by its perplexity. However, finding the prompt with the lowest perplexity is challenging, given the enormous space of possible prompting phrases. In this paper, we propose monotonic paraphrasing (MonoPara), an end-to-end decoding strategy that paraphrases given prompts or instructions into their lower perplexity counterparts based on an ensemble of a paraphrase LM for prompt (or instruction) rewriting, and a target LM (i.e. the prompt or instruction executor) that constrains the generation for lower perplexity. The ensemble decoding process can efficiently paraphrase the original prompt without altering its semantic meaning, while monotonically decrease the perplexity of each generation as calculated by the target LM. We explore in detail both greedy and search-based decoding as two alternative decoding schemes of MonoPara. Notably, MonoPara does not require any training and can monotonically lower the perplexity of the paraphrased prompt or instruction, leading to improved performance of zero-shot LM prompting as evaluated on a wide selection of tasks. In addition, MonoPara is also shown to effectively improve LMs’ generalization on perturbed and unseen task instructions.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in zero-shot decision making Gonen et al. (2023); Schick and Schütze (2021); Brown et al. (2020) and instruction following Jiang et al. (2023); Köpf et al. (2023); Touvron et al. (2023b); Taori et al. (2023); Chiang et al. (2023); Ouyang et al. (2022). However, there can be significant variance in the performance of seemingly similar prompts Zhao et al. (2021); Lu et al. (2022c); Webson and Pavlick (2022); Gonen et al. (2023). Despite efforts of studies on prompting LMs Shin et al. (2020); Li and Liang (2021); Gao et al. (2021); Ding et al. (2022); Sanh et al. (2021); Kojima et al. (2022), it is still challenging to develop high-quality prompts that can induce better performance for varying tasks on evolving models in an effort-saving manner.
@@
One consensus reached by recent studies is the inverse relationship between a prompt’s perplexity and its task performance Kumar et al. (2023); Gonen et al. (2023). This stems from the intuition that the frequency of a prompt (or an instruction) [^1] We use the terms “prompt” and “instruction” interchangeably in this paper as they both refer to a natural language command for zero-shot inference. appearing in the (pre-)training data positively influences the model’s familiarity with it, thereby enhancing its capability to perform the described task Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023). As a result, existing attempts use specific criteria, especially perplexity Gonen et al. (2023); Ye and Durrett (2023); Kumar et al. (2023), for selecting prompts from a collection of candidates. For example, Gonen et al. (2023) builds a large prompt pool for each task and selects the one with the lowest perplexity. However, the performance of such a method is limited by the size, quality, and even availability of the candidate pool. Searching the prompt with the lowest perplexity for a particular task remains challenging due to the vast expanse of potential prompts.
@@
To address this challenge, we propose a novel progressive prompt refining approach, namely monotonic paraphrasing (MonoPara), that can proactively rewrite the given prompt into its low-perplexity counterpart without compromising its expressiveness of the task. During prompt (or instruction) refinement, MonoPara conducts an ensemble decoding process of a paraphrasing model together with the target LM. The paraphrase model is instructed to rewrite the given prompt of a specific task, while the target model, which is later on used for task inference, provides a constraint that aims to lower the perplexity of the generation. MonoPara iteratively decodes tokens based on the ensemble of these two models. Intuitively, the paraphrase model can genuinely paraphrase the original prompt without altering its semantic meaning and remain instructive for the given task, while the target model restricts the search space of paraphrased tokens to those with low perplexity, resulting in a task prompt that the target model is more familiar with.
@@
Further, for the decoding process of MonoPara, we explore two decoding schemes. In addition to greedy decoding (Section 2.3) that iteratively generates the next token based on the weighted combination of predicted probabilities from both models, we also explored search-based decoding (Section 2.4). The search-based decoding scheme follows the look-ahead decoding paradigm Lu et al. (2022b), which keeps several sequence candidates and maintains the one with the lowest perplexity scored by the target model. Compared to recent prompt refinement approaches Gonen et al. (2023); Kumar et al. (2023); Shum et al. (2023), MonoPara needs zero training effort and far less computational cost, and do not require pre-existence of multiple prompt or instruction candidates Gonen et al. (2023). We find that by an ensemble of two LMs, we are able to consistently decode lower perplexity prompts, which would in turn bring about better generalization of LMs on downstream tasks.
@@
Our contributions are three-fold. First, we raise the intuition that the greedy generation of prompted paraphrasing is not monotonic in perplexity (Section 2.1). Second, we explore and compare two distinct decoding strategies that proactively refine prompts by monotonically lowering their perplexity and maintaining their expressiveness of the task. Third, we conduct experimentation for both prompt refinement for regular LMs and instruction refinement for instruction-tuned LMs to illustrate that monotonic paraphrasing improves the generalization of LM prompting.
@@ 2 Monotonic Paraphrasing
We propose MonoPara as an ensemble-based decoding method that can refine a given prompt by paraphrasing it into a low-perplexity counterpart. We first provide intuition of MonoPara in Section 2.1 and the basic of prompt paraphrasing in Section 2.2. Then, we formally define two decoding schemes of MonoPara in Section 2.3 and Section 2.4.
@@ 2.2 Prompt Paraphrasing
Specifically, we denote the input of paraphrase model, including the concatenation of the system prompt $x_{sys}$ and the original prompt $x_{ori}$ that the model is instructed to rewrite, as $[x_{sys},x_{ori}]=(x_{1},\dots,x_{n})$, where $x_{i}$ is a token in the vocabulary $V$. Suppose the decoder of a pre-trained autoregressive language model $P_{para}$ generates continuations of length $m$ as the requested paraphrase, denoted as $x_{para}=(x_{n+1},\dots,x_{n+m})$, based on the system prompt $x_{sys}$ and original prompt $x_{ori}$. At decoding time, $P_{para}$ iteratively decodes one token at a time by conditioning on the preceding context:
@@ 2.3 Ensemble-based (Greedy) MonoPara
By combining the prediction probabilities of both models, the ensemble can navigate the complex landscape of language generation more effectively and precisely. The paraphrase model $P_{para}$ keeps the rewritten prompt (or instruction) remain with the original intent, while the target model $P_{tar}$ evaluates the paraphrase’s likelihood or confidence, up-weighting any generation samples that lead towards low-perplexity outcomes. Based on the ensemble, the next token is selected by maximizing the weighted sum of logarithmic probabilities from both models, considering both the instruction and the generated text thus far.
@@ 2.4 Search-based MonoPara
To further enhance the efficiency of decoding prompts of lower perplexity, we define the search-based decoding strategy for MonoPara (as shown in the upper half of Figure 2). To take more steps of token generation into account and expand the search space, we leverage look-ahead decoding to search for the sequence that exhibits the lowest perplexity from an even broader candidate space. Following Lu et al. (2022a), each step of our search-based decoding consists of (i) expanding a set of candidate next-tokens, (ii) scoring each candidate, and (iii) selecting the $k$ best candidates (Algorithm 2):

### v3

@@ Abstract
Performance of large language models (LLMs) may vary with different prompts or instructions for even the same task. One commonly recognized factor for this phenomenon is the model’s familiarity with the given prompt or instruction, which is typically estimated by its perplexity. However, finding the prompt with the lowest perplexity is challenging, given the enormous space of possible prompting phrases. In this paper, we propose monotonic paraphrasing (MonoPara), an end-to-end decoding strategy that paraphrases given prompts or instructions into their lower perplexity counterparts based on an ensemble of a paraphrase LM for prompt (or instruction) rewriting, and a target LM (i.e. the prompt or instruction executor) that constrains the generation for lower perplexity. The ensemble decoding process can efficiently paraphrase the original prompt without altering its semantic meaning, while monotonically decreasing the perplexity of each generation as calculated by the target LM. We explore in detail both greedy and search-based decoding as two alternative decoding schemes of MonoPara. Notably, MonoPara does not require any training and can monotonically lower the perplexity of the paraphrased prompt or instruction, leading to improved performance of zero-shot LM prompting as evaluated on a wide selection of tasks. In addition, MonoPara is also shown to effectively improve LMs’ generalization on perturbed and unseen task instructions. [^1] Our code is available at https://github.com/luka-group/MonoPara.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in zero-shot decision making Gonen et al. (2023); Schick and Schütze (2021); Brown et al. (2020) and instruction following Jiang et al. (2023); Köpf et al. (2023); Touvron et al. (2023b); Taori et al. (2023); Chiang et al. (2023); Ouyang et al. (2022). However, there can be significant variance in the performance of seemingly similar prompts Zhao et al. (2021); Lu et al. (2022c); Webson and Pavlick (2022); Gonen et al. (2023); Yan et al. (2024). Despite efforts of studies on prompting LMs Shin et al. (2020); Li and Liang (2021); Gao et al. (2021); Ding et al. (2022); Sanh et al. (2021); Kojima et al. (2022), it is still challenging to develop high-quality prompts that can induce better performance for varying tasks on evolving models in an effort-saving manner.
@@
One consensus reached by recent studies is the inverse relationship between a prompt’s perplexity and its task performance Kumar et al. (2023); Gonen et al. (2023). This stems from the intuition that the frequency of a prompt (or an instruction) [^2] We use the terms “prompt” and “instruction” interchangeably in this paper as they both refer to a natural language command for zero-shot inference. appearing in the (pre-)training data positively influences the model’s familiarity with it, thereby enhancing its capability to perform the described task Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023). As a result, existing attempts use specific criteria, especially perplexity for selecting prompts from a collection of candidates. For example, Gonen et al. (2023) builds a large prompt pool for each task and selects the one with the lowest perplexity. However, the performance of such rewrite-then-select methods Jiang et al. (2020); Zhou et al. (2022); Gonen et al. (2023); Prasad et al. (2023) is limited by the size, quality, and even availability of the candidate pool. Searching the prompt with the lowest perplexity for a particular task remains challenging due to the vast expanse of potential prompts.
@@
To address this challenge, we propose a novel progressive end-to-end prompt refining approach, namely monotonic paraphrasing (MonoPara), that can proactively rewrite the given prompt into its low-perplexity counterpart without compromising its expressiveness of the task. During prompt (or instruction) refinement, MonoPara conducts an ensemble decoding process of a paraphrasing model together with the target LM. The paraphrase model is instructed to rewrite the given prompt of a specific task, while the target model, which is later on used for task inference, provides a constraint that aims to lower the perplexity of the generation. MonoPara iteratively decodes tokens based on the ensemble of these two models. Intuitively, the paraphrase model can genuinely paraphrase the original prompt without altering its semantic meaning and remain instructive for the given task, while the target model restricts the search space of paraphrased tokens to those with low perplexity, resulting in a task prompt that the target model is more familiar with.
@@
Further, for the decoding process of MonoPara, we explore two decoding schemes. In addition to greedy decoding (Section 2.3) that iteratively generates the next token based on the weighted combination of predicted probabilities from both models, we also explored search-based decoding (Section 2.4). The search-based decoding scheme follows the look-ahead decoding paradigm Lu et al. (2022b), which keeps several sequence candidates and maintains the one with the lowest perplexity scored by the target model. Compared to recent prompt refinement approaches Gonen et al. (2023); Kumar et al. (2023); Shum et al. (2023), MonoPara needs zero training effort and far less computational cost, and do not require the pre-existence of multiple prompt or instruction candidates Gonen et al. (2023). We find that by an ensemble of two LMs, we are able to consistently decode lower perplexity prompts, which would, in turn, bring about better generalization of LMs on downstream tasks.
@@
Our contributions are three-fold. First, we propose MonoPara, a prompt refinement method that paraphrases prompts to be more familiar with the target model for boosted generalization and task performance. Second, we explore and compare two distinct decoding strategies that proactively refine prompts by monotonically lowering their perplexity and maintaining their expressiveness of the task. Third, we conduct experimentation for both prompt refinement for regular LMs and instruction refinement for instruction-tuned LMs to illustrate that monotonic paraphrasing improves the generalization of LM prompting.
@@ 2 Monotonic Paraphrasing
We propose MonoPara as an ensemble-based decoding method that can refine a given prompt by paraphrasing it into a low-perplexity counterpart. We first provide the intuition of MonoPara in Section 2.1 and the basics of prompt paraphrasing in Section 2.2. Then, we formally define two decoding schemes of MonoPara in Section 2.3 and Section 2.4.
@@ 2.2 Prompt Paraphrasing
The System Prompt for Paraphrasing Generate ONE paraphrase of the following sentence.
Specifically, we denote the input of the paraphrase model, including the concatenation of the system prompt $x_{sys}$ and the original prompt $x_{ori}$ that the model is instructed to rewrite, as $[x_{sys},x_{ori}]=(x_{1},\dots,x_{n})$, where $x_{i}$ is a token in the vocabulary $V$. Suppose the decoder of a pre-trained autoregressive language model $P_{para}$ generates continuations of length $m$ as the requested paraphrase, denoted as $x_{para}=(x_{n+1},\dots,x_{n+m})$, based on the system prompt $x_{sys}$ and original prompt $x_{ori}$. At decoding time, $P_{para}$ iteratively decodes one token at a time by conditioning on the preceding context:
@@ 2.3 Ensemble-based (Greedy) MonoPara
By combining the prediction probabilities of both models, the ensemble can navigate the complex landscape of language generation more effectively and precisely. The paraphrase model $P_{para}$ keeps the rewritten prompt (or instruction) remains with the original intent, while the target model $P_{tar}$ evaluates the paraphrase’s likelihood or confidence, up-weighting any generation samples that lead towards low-perplexity outcomes. Based on the ensemble, the next token is selected by maximizing the weighted sum of logarithmic probabilities from both models, considering both the instruction and the generated text thus far.
@@ 2.4 Search-based MonoPara
To further enhance the efficiency of decoding prompts of lower perplexity, we define the search-based decoding strategy for MonoPara (as shown in the upper half of Figure 2). To take more steps of token generation into account and expand the search space, we leverage look-ahead decoding to search for the sequence that exhibits the lowest perplexity from an even broader candidate space. Following Lu et al. (2022a), each step of our search-based decoding consists of (i) expanding a set of candidate next-tokens, (ii) scoring each candidate, and (iii) selecting the $k$ best candidates (Algorithm 2 in Appendix A):
@@
3 Experiment
3.1 Task I: Prompt Refinement
Task Description
To inspect if monotonic paraphrasing can refine a prompt and elicit better performance on prompting an LLM as designed, we use Super-NaturalInstruction (Sup-NatInst for short, Wang et al., 2022). This benchmark consists of 1,616 diverse NLP tasks and their expert-written prompts for evaluating the zero-shot generalizability of LLMs on a variety of NLP tasks. In Sup-NatInst, each task is paired up with an instruction that consists of the task definition for mapping an input text to a task output and several examples for demonstrating the desired or undesired output. To make our evaluation more challenging, we only use the task definition as a prompt for the target model to perform the task. During test time, the target model is prompted with a concatenation of Sup-NatInst task description (or its paraphrase), test sample, and multiple choices of answers. Since Sup-NatInst provides only one description of definition for each task, we use GPT4 Achiam et al. (2023) to generate $4$ more Sup-NatInst-like task descriptions for each involved task, which results in $5$ prompts for each task in total.
Datasets
Following Gonen et al. (2023), we choose $7$ classification tasks from Huggingface Datasets  [^3] https://huggingface.co/docs/datasets/index to have a set of diverse tasks: (i) GLUE Cola Warstadt et al. (2019) for grammatical acceptability discrimination; (ii) Newspop Moniz and Torgo (2018) for news classification; (iii) AG News Zhang et al. (2015) for news classification; (iv) IMDB Maas et al. (2011) for movie review sentiment analysis; (v) DBpedia Lehmann et al. (2015) for topic classification; (vi) Emotion Saravia et al. (2018) for tweet emotion classification; and (vii) Tweet Offensive Barbieri et al. (2020) for offensive tweet discrimination. We sample 1,000 examples from each dataset for prompt evaluation. For these tasks, the prompt follows the input, and at the end of each prompt, we add the choices of classes following Gonen et al. (2023): “Choices: X, Y, Z. Answer: ” as it helps with accuracy.
Evaluation Metrics
To inspect whether monotonic paraphrasing can refine a prompt into its low-perplexity counterpart and further enhance the generalization of LLM prompting, we use PPL (perplexity), Acc (accuracy), and BS (BERTScore, Zhang* et al., 2020) for evaluation. (1) Perplexity. Following Gonen et al. (2023), we define the perplexity of the prompt as the perplexity of the full prompt sequence, including the input itself and the choices of labels. To avoid noise when computing perplexity, the prompts are instantiated with 1,000 random examples of the dataset (we keep the same selected examples for performance evaluation), and the perplexity is averaged across all instantiated prompts. The correlation between the perplexity of a standalone prompt and the average perplexity over instantiated ones is illustrated in Appendix E. (2) BERTScore. As a paraphrasing scheme, MonoPara is also evaluated by its performance in semantic alignment with the source sentence based on pre-trained BERT Devlin et al. (2019). (3) Accuracy. To compute accuracy, for each test sample, we obtain the model’s predicted probability of all classes and choose the highest ranking class as the prediction of the model. The accuracy is calculated over the 1,000 samples for each task.
Mistral 7B
Acc 43.42 ± 1.16 81.08 ± 7.2 73.88 ± 3.18 86.00 ± 0.81 71.02 ± 1.70 46.80 ± 1.12 65.06 ± 1.04
SPELL PPL 8.24 ± 0.88 15.05 ± 4.15 15.99 ± 3.77 15.76 ± 1.42 25.36 ± 6.80 15.67 ± 5.21 14.31 ± 0.93
Para BS $8$ 0.74 87.47 ± 0.84 86.32 ± 2.08 90.23 ± 0.99 86.96 ± 0.93 90.08 ± 0.35 89.93 ± 0.67
Starling 7B
Acc 46.46 ± 6.70 89.84 ± 0.91 83.72 ± 0.26 92.84 ± 2.7 92.02 ± 0.64 54.82 ± 1.81 64.20 ± 0.17
SPELL PPL 44.96 ± 18.23 54.60 ± 18.72 16.81 ± 0.91 56.34 ± 15.67 28.66 ± 9.47 32.79 ± 8.53 33.56 ± 3.83
Para BS 90.35 ± 0.67 88.35 ± 0.51 88.46 ± 0.98 90.04 ± 1.04 87.06 ± 0.79 90.38 ± 0.45 79.93 ± 11.52
Table 1: Results of Task I (Prompt Refinement). Ori. refers to the original prompts. Para refers to paraphrasing the original prompts without perplexity constraint. Mono-E refers to ensemble-based MonoPara. Mono-S refers to search-based MonoPara. The best accuracy for each task is in bold, while the lowest perplexity and the highest BERTScore for each task is in blue and orange respectively. The coefficient $\alpha$ is fixed as $0.5$ for all the results.
Models
We study two auto-regressive models, Mistral-7B [^4] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023)and Starling-7B [^5] https://huggingface.co/berkeley-nest/Starling-LM-7B-alpha Zhu et al. (2023a), for their significant ability in instruction following and paraphrasing. In this paper, we keep the paraphrase model the same as the target model. Please refer to Appendix B for the results on Llama-3-8B [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B AI@Meta (2024).
Baselines
We compare our method with two baselines: SPELL Gonen et al. (2023) and vanilla LLM-based paraphrasing (Para). Based on similar intuition that low perplexity and better performance exhibit strong correlation, SPELL (Selecting Prompts by Estimating LM Likelihood) ranks and selects the prompts with the lowest perplexity for a given task after creating a set of prompt candidates manually and expanding them to hundred-scale using automatic paraphrasing and back-translation. We use the prompt candidates provided by Gonen et al. (2023) and rank them by the perplexity w.r.t. the target model. Para paraphrases source prompts with greedy decoding (top-$1$ sampling) by directly prompting the target LLM with the instruction introduced in Section 2.2. Besides, we also implement reranking based on the prompt candidates generated by MonoPara for a fair comparison with SPELL. The implementation details and results are discussed in Appendix D.
Results
As shown in Table 1, the two variants of MonoPara successfully refine prompts to those with perplexities lower than the original ones, thereby enhancing model performance on 10 out of 14 scenarios (2 models, each on 7 datasets). Specifically, Mono-E (ensembled-based MonoPara) improves the accuracy of Mistral-7B by an average of 2.64% and that of Starling-7B by an average of 1.85% on seven different datasets. In contrast, directly paraphrasing the prompts without perplexity constraint (Para) results in prompts with higher perplexities than the original ones in 13 out of 14 scenarios, most of which lead to worse task performance. This highlights the importance of incorporating perplexity constraints when paraphrasing prompts. The correlation between task accuracy and prompt perplexity is analyzed in Appendix C.
Besides, our method outperforms SPELL with consistently lower perplexity and superior task performance with the refined prompt. This superiority stems from the fact that, although SPELL explores a vast search space with hundred-scale prompt candidates, the candidates are searched in an ad hoc manner without an optimization goal, leading to inefficiencies in prompt rewriting or paraphrasing. In contrast, our method incorporates the perplexity constraint that leads to optimized target-model perplexity during one single refined prompt generation process, leading to both enhanced task performance and refinement efficiency.
Moreover, MonoPara achieves superior performance in terms of both average prompt perplexity and BERTScore, indicating the synergistic role played by both the target model and the paraphrase model in prompt refinement. During decoding, the target model prioritizes achieving low perplexity, while the paraphrase model preserves the original prompt intent, resulting in a high BERTScore. We provide a detailed case study on BERTScore of Para and MonoPara in Appendix G to explain why MonoPara outperforms Para on BERTScore. Specifically, Mono-E outperforms Mono-S in some cases since Mono-S involves the additional hyperparameter of beam size which we did not tune. Additionally, having additional token options at each decoding step may introduce more randomness, potentially affecting the overall performance. That being said, Mono-S mostly exhibits a higher BERTScore than the simpler method Mono-E, indicating better quality of paraphrasing.
3.2 Task II: Robustness Evaluation on Instruction Perturbation
Task Description
PromptBench introduces perturbation to task instructions of a diverse set of tasks, such as sentiment analysis, grammar correctness, duplicate sentence detection, and natural language inference. It includes character-level perturbations using DeepWordBug Gao et al. (2018) to introduce typos, word-level using TextFooler Jin et al. (2020) to replace words with contextually similar words, and semantic-level by using prompts following the linguistic behavior of different languages. For semantic-level perturbations, the adversary constructs prompts using various languages, such as Chinese, French, Arabic, Spanish, Japanese, and Korean, and then translates these prompts into English. By exploiting the nuances and idiosyncrasies of different languages during translation, it can introduce subtle ambiguities, grammatical errors, or inconsistencies in the input prompt. The perturbed prompt itself is still in English.
Figure 3: Model’s average accuracy across $4$ GLUE datasets, with each dataset having six instructions with perturbation added at character, word, and semantic levels. Mono-E has consistent improvement in accuracy across all types of perturbation compared to vanilla paraphrasing.
Datasets
We select $4$ GLUE tasks Wang et al. (2018) from Huggingface Datasets for evaluation: (i) Cola, which is also used for Task I; (ii) SST-2 Socher et al. (2013) for two-way sentiment classification; (iii) MNLI Wang et al. (2018) for multi-genre inference identification; and (iv) MRPC Dolan and Brockett (2005) for paraphrase identification. We adopt the same instruction format as that of Task I.
Model
For this task, we use Alpaca [^7] https://huggingface.co/tloen/alpaca-lora-7b. We use a different language model for this task to demonstrate the robustness and versatility of our proposed method across various models. Taori et al. (2023), a model instruction-tuned from the LLaMA model Touvron et al. (2023a) on the 52k self-instruct dataset. The paraphrase model is also the same as the target model.
Results
As shown in Figure 3, we compare the performance of original prompts, prompts paraphrased without perplexity constraint, and our method Mono-E. Across all kinds of instruction perturbations, our method achieves consistent accuracy improvement by up to 10.13%. This notable increase in performance demonstrates that MonoPara not only refines prompts for improved task performance but also enhances model robustness against various instruction perturbations. We also show the model performance with unperturbed prompts and compare the degradation from perturbation in Appendix F to show the robustness of Mono-E.
Content BertScore PPL
Ori Instruction
Read the provided excerpt and choose between
‘positive’ and ‘negative’ to describe its sentiment Iyk1lJt8yw:
- 78.12
Para Determine the sentiment of the given excerpt as either positive or negative 81.26 44.29
[…]
3.3 Effect of Coefficient
We further analyze the effect of coefficient $\alpha$ between the target model and the paraphrase model in MonoPara. Table 2 presents a case study of $\alpha$ on the ensemble-based decoding variant (Mono-E) with a prompt under character-level perturbation. Obviously, a larger value of $\alpha$ amplifies the influence of the target model, resulting in a stronger perplexity constraint and monotonically lower perplexity. Consequently, the impact of the paraphrase model is diminished, leading to a lower BERTScore. By adjusting $\alpha$, the user can find a balanced point to achieve better prompts that strike a balance between semantic fidelity to the original prompt intent and linguistic familiarity to the target model. Detailed results and analysis are presented in Appendix H and Table 9.
4 Related Work
Different from existing paraphrasing techniques, this work conducts constrained decoding methods that seek to monotonically decrease the perplexity in the process of paraphrasing. This innovative approach serves as a general approach to rewrite and refine prompts or task instructions towards their more familiar counterparts for LMs, seeking to enhance the generalization of LM prompting without the need for any training effort.
Prompt Refinement. Prompt refinement has been actively explored in recent years with the aim of selecting, retrieving, or even generating prompts that lead to improved zero-shot performance of LM prompting. When an end-task objective is known, prior works on prompt refinement approaches often set the end-task performance as the objective, and leverage approaches such as gradient-based search, similar to continuous prompts but with projections onto a discrete vocabulary Shin et al. (2020). Other works have explored edit-based enumeration Prasad et al. (2023), reinforcement learning Deng et al. (2022), and large language model continuation and filtering Zhou et al. (2022). As for general-purpose LLMs, a dedicated prompt refinement approach may not assume the pre-existence of any specific end tasks. Hence, more recent works Iter et al. (2023); Gonen et al. (2023); Wang et al. (2023); Lee et al. (2023) explored the heuristic criterion based on the familiarity of the language model with the language contained in the prompt, as measured by its perplexity. Lower perplexity prompts are preferred as they are expected to perform better across a wide range of tasks.
5 Conclusion
In this paper, we propose monotonic paraphrasing (MonoPara) for automatically and iteratively generating low-perplexity prompts for given LLMs, which can lead to better task performance. Following the high-level idea of paraphrasing prompts with perplexity constraints from the target LLM, we design ensemble-based and search-based decoding strategies for efficient prompt refinement. Experiments on prompt variations of unseen tasks and instructions demonstrate the effectiveness of MonoPara in reducing prompt perplexity while enhancing task performance and model robustness. Future work may apply MonoPara on other scenarios, such as searching stealthy prompts for red-teaming. Replacing perplexity with other criteria, such as complexity Li et al. (2024), for different purposes of constrained generation is also a meaningful research direction.
Acknowledgement
We thank the anonymous reviewers for their valuable comments. Qin Liu was supported by a departmental fellowship. Fei Wang was supported by the Amazon ML Fellowship. Tianyi Yan was supported by the CURVE Fellowship. Muhao Chen was supported by the DARPA FoundSci Grant HR00112490370, the NSF of the United States Grant ITE 2333736, and an Amazon Research Award.
Limitations
The current investigation of MonoPara has the following limitations. First, using the target model as the paraphrase model is a natural technical choice. However, examining other paraphrasing models, especially those customized for paraphrasing, may lead to better performance. Second, our proposed method can not be applied to black-box LLMs such as ChatGPT [^8] https://chat.openai.com, PaLM Chowdhery et al. (2023), Claude [^9] https://www.anthropic.com/claude, etc. This limitation arises from the necessity of access to the decoding phase, which in turn comes with far less computational cost due to zero training, and does not require the pre-existence of multiple prompt or instruction candidates. Third, while we conduct experiments on a wide range of tasks, their original prompts and responses are relatively short. This actually limits the operational space of MonoPara. Experiments on tasks with longer inputs and outputs may provide additional evidence of MonoPara’s effectiveness.
Appendices
Appendix A Algorithm for Search-based MonoPara